OpenStudy (toman):

What is the area of a circle whose circumference is 1447 meters? (Use π = 3.14)

OpenStudy (anonymous):

Do you know the formulas for area and circumference?

OpenStudy (toman):

No.

OpenStudy (anonymous):

A = πr² and C = 2πr

OpenStudy (toman):

Oh okay.

OpenStudy (toman):

I have to find half of 1447, so I can plug it into the "r"

OpenStudy (anonymous):

You can use the circumference formula to find r. Then use the r to get area.

OpenStudy (anonymous):

So not just half of 1447. You also have to divide by 3.14

OpenStudy (toman):

Yes. Okay.

OpenStudy (toman):

Since it's \(2πr\), wouldn't it be \(2 \times\ 3.14 \times\ r = 1447\) ??

OpenStudy (anonymous):

yes that's right

OpenStudy (toman):

So now it's \(6.28 \times\ r\) ??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

That's what I got

OpenStudy (toman):

\(~r=\frac{ 1447 }{ 6.28 } = 230.41\)

OpenStudy (toman):

I meant that..

OpenStudy (toman):

So now I have the radius.

OpenStudy (anonymous):

So now use that to find A

OpenStudy (toman):

Okay.

OpenStudy (toman):

\(A=3.14 \times\ 230.41 \times\ 230.41\)

OpenStudy (anonymous):

yep

OpenStudy (wolf1728):

Did you multiply it out yet?

OpenStudy (toman):

I got \(166698.73\). I'm not sure if that's right.

OpenStudy (anonymous):

Yes I think that's a good answer. I didn't round and I got 166,704.54. I think it's an acceptable amount of error for such a large number.

OpenStudy (toman):

Okay. Thank you so much.

OpenStudy (anonymous):

you're welcome

OpenStudy (wolf1728):

Heck I got 166,704.54 also wow this openstudy has bounced me out 2 times already.

OpenStudy (toman):

It's okay. Thank you, as well. c: